OpenStudy (anonymous):

N(20)=4000(2)^(20/40) solve for N, round to nearest whole number

OpenStudy (anonymous):

Rewritten:\[20N=4000\sqrt{2}\]The 20/40 reduced to 1/2 and something being taken to the power of one half is the same as taking the square root of it.

OpenStudy (anonymous):

I've tried solving that in that format to find my number, I can solve it for other numbers but with the 20 in it, it's wrong every stinkin time...

OpenStudy (anonymous):

Unless the 4000 is in the square root with the 2 in which case you would get\[20N=\sqrt{8000}\]

OpenStudy (anonymous):

With the first way, N=283, and the second, N=4. Either of those work?

OpenStudy (anonymous):

I got those answers, nope.. not right. at this point I can't even give more guesses. thanks for helping. I can now solve for N(10)=4000(2)^(10/20)